# LLM-Consensus: Multi-Agent Debate for Visual Misinformation Detection

### v2

@@ 4.2 Experimental Setup
All experiments were run on 8 A40 (46GB) Nvidia GPU server. The estimated cost of processing one data sample using LLM-Consensus is \$0.24. Inference times range from 5 to 15 seconds.
Debate Setup:
We conduct experiments to select the best debating configuration using the LLaVA model (Liu et al., 2024b). The experiments are carried out on a smaller subset containing 1000 test samples of the main NewsCLIPpings test dataset. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
External Retrieval Module:
We use the Bing Visual Search API (vis, ) to run an image-based reverse search. Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library (new, ). Finally, we use Llama-13B (Touvron et al., 2023) to summarise the text obtained from the top $k=3$ web pages. This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
Baselines and Prior Work:
We compare LLM-Consensus to existing pretrained multi-modal baselines including CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2024b). We also compare performance against GPT-4o (OpenAI & et al., 2024; OpenAI, ). The models are presented with the image and caption pair and asked if the pair is misinformation. The models are further prompted to explain their reasoning. We also show results for two baseline methods trained from scratch, namely EANN (Wang et al., 2018) and SAFE (Zhou et al., 2020). We further compare LLM-Consensus to DT-Transformer (Papadopoulos et al., 2023), CCN (Abdelnabi et al., 2022), Sniffer (Qi et al., 2024), VINVL (Huang et al., 2024), SSDL (Mu et al., 2023) and Neuro-Sym (Zhu et al., 2022).
@@ 4.3 Results
We present results for the experiments conducted to select the best debate setup as well compare the performance of LLM-Consensus against existing methods. We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
@@ 4.3.1 Comparing Debate Setups
We compare multiple debating setups using the LLaVA model, to select the best one for comparison with other works and further experimentation.
@@
Actor-Skeptic 69.5 66.1 69.4
Judged Debate 66.7 66.7 61.5
Debate with Disambiguation 77.8 74.7 82.6
Table 1: Performance comparison between different debate setups: The Async_Debatehuman where the model has external context and believes it is debating a human being is the best setup.
Table 1 shows that the Async_Debatehuman setup where the agent has access to external information performs the best of all the debating configurations. We also report results for the Async_Debatehuman setup without access to external information to emphasise the importance of external information for the problem of misinformation detection in the news domain. The external retrieval of information significantly boosts performance. We also observe a significant performance increase when the agent believes it is conversing with a human instead of another AI agent. Qualitatively, the agent considers the other agent’s responses more critically and with more seriousness when it believes that the agent is a human. Further, the asynchronous debate setup benefits from the ensemble of agents, which is not present in the actor-skeptic setup, where only one agent is responsible for generating the responses. The generation of disambiguation queries within the same response, confuses the agents and even deviates them from their own chain of thought. We believe this accounts for the counter-intuitively performance of this method where agents perform worse with more information. The judged debate setup focuses on enforcing agents to structure their responses in a way that will convince a judge. The agents also debate with opposite stances and do not have the option of changing their stance mid-debate. This can further confuse the judge and lead to incorrect decisions. This is resolved in the Async_Debatehuman set up where agents are given complete freedom over their initial opinions, as well as their opinions during the debate. If they believe they are convinced by the other agents’ arguments, they can choose to change their response and the debate ends. Based on the results from Table 1, we choose the best-performing debate set-up, i.e. Async_Debatehuman (with external information) as the debate configuration for further experimentation and comparison.
@@ 4.3.2 Performance Comparison
We present our results on the NewsCLIPpings dataset against existing out-of-context detection methods discussed in section 4.2.
@@
LLM-Consensus (ours) 90.8
Table 2: Performance comparison between our model and baselines: LLM-Consensus (with GPT-4o) out performs all related work. Note: the GPT-4o# setup is identical to LLM-Consensus with the absence of a multi-agent debate, here only a single agent with access to external information is considered.
Table 2 shows the comparison between our system and existing methods. We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o (OpenAI & et al., 2024; OpenAI, ) model. Sniffer (Qi et al., 2024), being the only work comparable in performance to ours, is finetuned extensively to adapt it to the NewsCLIPpings dataset. While we do not provide a quantitative assessment of explanations by LLM-Consensus, we do believe our system produces more coherent, detailed and comprehensive explanations when compared to other baselines. This is attributed to the fact that in a multi-agent setup, we have multiple context windows which leads to more coherent and relevant final explanations. We leave the detailed analysis of these explanations and the development of the associated metrics as future work. We also note that the debate paradigm in itself is essential to the system performance. We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model.
We also note that single multi-modal models, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work. This can be attributed to the necessity for external context for misinformation detection in the news domain and the lack of diverse perspectives that arise naturally in a multi-agent framework. Therefore, these standalone models, while promising, currently are unable to detect misinformation effectively. These models require additional integration into more comprehensive pipelines, as done in this work. In line with previous work, we also note that baselines trained from scratch, such as SAFE (Zhou et al., 2020) and EANN (Wang et al., 2018) perform worse than pretrained multi-modal models. This further concretizes the fact that image-based OOC detection in the news domain requires strong world knowledge as well as advanced multi-modal reasoning capabilities.
4.3.3 Ablations
To analyze the importance of each component of the LLM-Consensus framework, we conduct ablation experiments. Specifically, we evaluate the effect of using LLaVA against GPT-4o, the impact of the external retrieval module, and the power of the multi-agent debate framework.
@@ 4.3.3 Ablations
Table 3: Ablation: Quantitative evaluation of each component of LLM-Consensus on classification performance.
@@
We observe that the combination of GPT-4o, the external retrieval module, and the multi-agent debate framework yields the highest performance across all metrics, with 90.8% accuracy, 85.5% precision, and 99.0% recall, demonstrating the value of combining these components. The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7 (as seen in Table 2) to 90.2, underscoring its role in enabling contextual reasoning and refining predictions. Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3 to 85.5 and recall from 90.1 to 99.0. Meanwhile, retrieval contributes more substantially to LLaVA’s performance gains, likely due to GPT-4o’s broader world knowledge. Without external retrieval or the debate framework, the performance drops, emphasizing the critical role of these components in achieving state-of-the-art results.
5 User Study
We conducted a user study to evaluate the effectiveness of our system in detecting and explaining misinformation. While model performance in misinformation detection is easy to quantify, there are no established metrics for assessing the quality of its explanations, making the user study essential. Participants were grouped into three categories based on their professions: Journalists, AI Academics, and Others (see Appendix A.4 for details).
During the study, participants reviewed ten image-text pairs and decided whether each pair constituted misinformation. They also rated their confidence on a scale of 0 to 10. After submitting their initial answers, participants were shown LLM-Consensus’s AI-generated explanations and asked to reconsider their responses. As shown in Table 4, the system outperformed the average human performance in both cases, with and without AI insights. These results demonstrate LLM-Consensus’s potential as a reliable assistive tool for OSINT research, capable of detecting and explaining misinformation with minimal human intervention.
[…]
Humans $60.3\pm 13.5$
Humans+LLM-Consensus $76.7\pm 12.2$
LLM-Consensus $\mathbf{80.0\pm 0.0}$
Table 4: Performance comparison between different study setups: LLM-Consensus outperforms humans with and without AI assistance.
Group-wise analysis, shown in Table 5, reveals significant performance gains across all groups, with results approaching those of professional journalists. Confidence levels (out of 10) are comparable across groups and generally increase after using LLM-Consensus insights. Thus, LLM-Consensus can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
Metric Journalists Academics Others
Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
Accuracy (with LLM-Consensus) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
Confidence (with LLM-Consensus) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
Table 5: Performance comparison: LLM-Consensus improves performance across all participant groups.
6 Conclusion and Future Work
Misinformation detection has become a pressing issue in recent times. With the ever-advancing capabilities of vision and language models, the detection of OOC image use has become a very difficult task. In this work, we explore the question of whether it is possible for multiple AI agents to pool their contextual knowledge and converge to a common prediction in order to identify instances of misinformation. We identify Asynchronous_Debatehuman as the most optimal communication setup for AI models. We observe significant performance improvement when the models believe they are debating against a human instead of another AI agent. We observe that in this setup, models tend to be more involved and open to changing their opinions. Our method also allows for agents to have freedom of opinion which they may change mid-debate. Agents in such a setting show enhanced abilities to critically evaluate an argument and pick up on minute inconsistencies.
Our final system, LLM-Consensus, achieves state-of-the-art performance in misinformation detection while offering clear, detailed explanations, thanks to our advanced external retrieval module. As a result, we observe significant improvements in out-of-context (OOC) detection for both experts and non-experts.
We identify several promising directions for future research. The community would benefit from a continuously updated benchmark dataset incorporating recent news articles and subtler inconsistencies. We plan to extend our methods to video-text pairs and multi-lingual content. We also believe that integrating more advanced models in the summarization pipeline will further enhance performance and are interested in comparing LLM-Consensus with multi-agent systems using external retrieval.
Finally, we conducted extensive user studies but recognize the importance of large-scale deployment in professional settings and the citizen intelligence community. This will provide valuable real-world feedback and uncover further opportunities for improvement. For a discussion of limitations, see Appendix A.1.
Appendix A Appendix
A.1 Limitations
Despite the strong performance of LLM-Consensus, several limitations remain. First, while our model excels at detecting out-of-context image-text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve. Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration. Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
Another important limitation is the potential risk that open-sourcing LLM-Consensus might allow adversaries to train models specifically designed to counter or evade detection by our system. As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness. However, despite these risks, we believe that open-sourcing remains the right path forward. Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining LLM-Consensus’s effectiveness in the long term. The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of LLM-Consensus.
A.2 Sample Image-Caption Pair in the News Domain
A.3 Prompts for LLM-Consensus
This is a summary of news articles related to the image: {}
CAPTION: {}
Note that the image is real. It has not been digitally altered.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 5: Initial prompt for independent opinion formation and response generation
This is what I think: {}.
Do you agree with me?
If you think I am wrong then convince me why you are correct.
Do not describe the image.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 6: Prompt for Debate Round 1
I see what you mean and this is what I think: {}.
Do you agree with me?
I want you to help me improve my argument and explanation.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 7: Prompt for Debate after Round 1
A.4 User Study

### v3

@@ 4.2 Experimental Setup
All experiments were run on 8 A40 (46GB) Nvidia GPU server. The estimated cost of processing one data sample using MAD-Sherlock with a GPT-4o backbone is \$0.24. Inference times range from 5 to 15 seconds.
Debate Setup
We conduct experiments to select the best debating configuration using the LLaVA model [23]. All experiments are run for $k=3$ rounds or until the agents converge (whichever is earlier).
External Retrieval Module
We use the Bing Visual Search API to run an image-based reverse search. Using the API we select the top $k=3$ pages in which the image appears and scrape the text from them using the Newspaper3k library. Finally, we use Llama-13B [39] to summarise the text obtained from the top $k=3$ web pages. This step is crucial since the web pages are usually news articles which contain large amounts of text which, when scraped and passed directly to the model, can exceed its maximum token length.
Baselines and Prior Work
The models are presented with the image and caption pair and asked if the pair is misinformation. The models are further prompted to explain their reasoning. We present comparisons to other explainable methods, specifically Sniffer [33], CRAVE [10], and MMD-Agent [24] in the following section. We also compare MAD-Sherlock to existing pretrained multi-modal baselines including CLIP [34], VisualBERT [19], InstructBLIP [9], LLaVA [23], and GPT-4o [30, 29] on the NewsCLIPpings dataset in Appendix A.4. We show results for two baseline methods trained from scratch on the NewsCLIPpings, namely EANN [41] and SAFE [44], in Appendix A.4. We further compare MAD-Sherlock to DT-Transformer [31], CCN [1], VINVL [17], SSDL [28], and Neuro-Sym [45] in Appendix A.4.
@@ 4.3 Results
We present results for the experiments conducted to select the best debate setup as well compare the performance of MAD-Sherlock against existing methods. We use classification accuracy as the primary performance metric for comparison based on quantitative analysis.
@@ 4.3.1 Comparing Debate Setups
We compare multiple debating setups using the LLaVA-v1.6-34B model, to select the best one for comparison with other works and further experimentation.
@@
Table 1: Performance comparison between different debate setups: We observe best performance with the The Async_Debatehuman set-up, in which the model has external context and believes it is debating a human.
Table 1 shows that Async_Debatehuman with external information outperforms all other debate configurations. To highlight the role of external retrieval, we also report results without it and observe a significant performance drop.
We find that agents perform better when they believe they are debating a human, engaging more critically with peer responses. The asynchronous setup also benefits from ensemble effects, unlike the actor-skeptic setup where only one agent generates responses.
In judged debates, agents must adhere to fixed stances and persuade a judge, which can lead to confusion and errors. In contrast, Async_Debatehuman allows agents to revise their views mid-debate, leading to clearer outcomes. Based on these findings, we adopt this setup—with external information—for all further experiments.
@@ 4.3.2 Performance Comparison
We present our results on the NewsCLIPpings, VERITE, and MMFakeBench datasets against existing out-of-context detection methods discussed in Section 4.2. We include additional comparisons against legacy methods in Appendix A.4.
Dataset Model Performance
Accuracy Precision Recall
NewsCLIPpings SNIFFER [33] 88.4 91.8 86.9
CRAVE [10] 85.0 85.0 85.0
MAD-Sherlock (ours) 90.8 85.5 99.0
VERITE CRAVE [10] 82.0 82.0 82.0
MAD-Sherlock (ours) 85.2 83.6 96.0
MMFakeBench MMD-Agent [24] 62.1 67.8 59.3
CRAVE [10] 78.0 83.6 69.7
MAD-Sherlock (ours) 83.8 87.2 90.1
Table 2: Performance comparison of LLM-Consensus against other explainable methods on various datasets: Performance of LLM-Consensus and baselines across NewsCLIPpings, VERITE, and MMFakeBench. Our model shows consistent generalization and superior performance across datasets.
Table 2 shows the comparison between our method and other explainable methods on the NewsCLIPpings, VERITE, and MMFakeBench datasets. We report state-of-the-art performance when using our proposed debate configuration with the GPT-4o [30, 29] model across all datasets. As a reminder, Sniffer [33] is finetuned extensively to adapt it to the NewsCLIPpings dataset. MMD-Agent represents the single-agent framework proposed by the authors of the MMFakeBench dataset [24]. The CRAVE framework introduced by Dey et al. [10] represents the most recent work across all three datasets, but our framework shows markedly superior performance across all datasets.
We find that our system produces coherent, detailed and comprehensive explanations. We also note that the debate paradigm in itself is essential to the system performance. We observe a drop in performance and quality of explanations when using an identical system configuration but with a single model. We believe that this follows from work by Mireshghallah et al. [27], which introduces and demonstrates the importance of contextual privacy. Introducing multiple context windows allows each agent to maintain its own role and informational perspective without leakage, which is not possible in single-agent systems.
We provide a qualitative sample of our framework’s explanations in Appendix A.5. We provide ablation studies on each component of the MAD-Sherlock pipeline in Appendix A.6.
We also note that single multi-modal models that do not do retrieval, including VisualBERT, CLIP, InstructBLIP, LLaVA and GPT-4o do not perform at par with other related work. We demonstrate these results in Appendix A.4. This can be attributed to the necessity for external context for misinformation detection and the lack of diverse perspectives that occur naturally in a multi-agent framework. These models require additional integration into more comprehensive pipelines, as done in this work.
5 User Study
We conducted a user study to assess our system’s ability to detect and explain misinformation—particularly important given the lack of standard metrics for evaluating explanation quality. Participants were grouped by profession: Journalists, Academics, and Others (see Appendix A.7).
Each participant reviewed ten image-text pairs, judged their veracity, and rated their confidence (0–10). After submitting initial responses, they viewed MAD-Sherlock’s explanations and revised their answers. As shown in Table 3, MAD-Sherlock outperformed average human accuracy both with and without AI assistance, highlighting its potential for improving public safety and trust.
Study Setup Average Accuracy
Humans $60.3\pm 13.5$
Humans+MAD-Sherlock $76.7\pm 12.2$
MAD-Sherlock $\mathbf{80.0\pm 0.0}$
Table 3: Performance comparison between different study setups: MAD-Sherlock outperforms humans with and without AI assistance.
Group-wise analysis, shown in Table 4, reveals significant performance gains across all groups, with results approaching those of professional journalists. Confidence levels (out of 10) are comparable across groups and generally increase after using MAD-Sherlock insights. Thus, MAD-Sherlock can substantially boost non-expert performance, making it valuable for citizen intelligence applications.
Metric Journalists Academics Others
Accuracy (only human) $70.0\pm 1.4$ $60.7\pm 1.4$ $56.7\pm 1.5$
Confidence (only human) $4.3\pm 2.1$ $3.2\pm 0.8$ $3.9\pm 1.2$
Accuracy (with MAD-Sherlock) $82.2\pm 0.9$ $79.3\pm 1.3$ $71.7\pm 1.1$
Confidence (with MAD-Sherlock) $5.3\pm 1.3$ $5.8\pm 1.4$ $5.8\pm 1.4$
Table 4: Performance comparison: MAD-Sherlock improves performance across all participant groups.
6 Conclusion and Future Work
Out-of-context (OOC) image misuse is an increasing challenge for misinformation detection, especially as vision-language models grow more powerful yet less interpretable. We explore whether multiple AI agents can collaboratively reason about context to improve prediction accuracy. Our strongest results come from the Asynchronous_Debatehuman setup, where agents believe they are debating a human. This setting promotes engagement, mid-debate revision, and better identification of subtle inconsistencies.
Our final system, MAD-Sherlock, achieves state-of-the-art performance while offering interpretable, evidence-based explanations—enabled by our external retrieval module. We observe substantial gains in OOC detection across both expert and non-expert users.
We see several avenues for future work, including a continuously updated benchmark with recent news and more nuanced inconsistencies, extending to video-text and multilingual inputs, and large-scale deployments in professional and citizen intelligence settings. We hope to engage with the community to better understand how agentic workflows can enhance online safety. For limitations, see Appendix A.1.
Appendix A Appendix
A.1 Limitations
Despite the strong performance of MAD-Sherlock, several limitations remain. First, while our model excels at detecting out-of-context image–text pairs, its reliance on external retrieval can lead to reduced accuracy when relevant context is unavailable or difficult to retrieve. Moreover, our framework cannot independently verify the factual correctness of externally retrieved news articles; the truthfulness of any source may be debated, potentially introducing misinformation into the reasoning process. Nevertheless, we observe that the model’s judgments align closely with the human-created labels used in these widely accepted datasets, underscoring its practical utility. Second, the quality of explanations is constrained to textual outputs, limiting multi-modal explanation capabilities such as image or video integration. Third, the system’s performance is sensitive to hyperparameter tuning, including the number of debate rounds and agents, which may require further optimization for broader use cases.
Another important limitation is the potential risk that open-sourcing MAD-Sherlock might allow adversaries to train models specifically designed to counter or evade detection by our system. As adversarial actors gain access to the source code, they could exploit its known strengths and weaknesses to develop countermeasures that diminish its effectiveness. However, despite these risks, we believe that open-sourcing remains the right path forward. Open-sourcing encourages transparency, collaboration, and rapid innovation, enabling the broader community to contribute improvements, detect vulnerabilities, and build on the system.
Moreover, by engaging the community, we can foster the development of more resilient and adaptive models that evolve in response to emerging adversarial techniques, thus maintaining MAD-Sherlock’s effectiveness in the long term. The collective strength of a diverse, open-source community can outweigh the potential threats posed by adversarial exploitation.
Future work will need to address these limitations to enhance the practical utility, robustness, and long-term resilience of MAD-Sherlock.
A.2 Sample Image-Caption Pair in the News Domain
A.3 Prompts for MAD-Sherlock
You are a misinformation detection expert in the news domain. You will look at image-caption pairs and decide if the given image is rightly used in the given news context. To further assist you, a summary of news articles related to the image will be provided. Based on this, you need to decide if the caption belongs to the image or if it is being used to spread false information to mislead people. Note that the image is real. It has not been digitally altered. Carefully examine the image for any known entities, people, watermarks, dates, landmarks, flags, text, logos and other details which could give you important information to better explain your answer. Remember in news articles images and captions are often related contextually and the caption need not exactly describe the image. The goal is to consider the contextual relationship between the image and caption based on the news articles and correctly identify if the image caption pair is misinformation or not and to explain your answer in detail. Think step by step and plan a detailed explanation for your answer.
Figure 5: Model initialization prompt.
This is a summary of news articles related to the image: {}
CAPTION: {}
Note that the image is real. It has not been digitally altered.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 6: Initial prompt for independent opinion formation and response generation with GPT-4o.
This is what I think: {}.
Do you agree with me?
If you think I am wrong then convince me why you are correct.
Do not describe the image.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 7: Prompt for first round of debate.
I see what you mean and this is what I think: {}.
Do you agree with me?
I want you to help me improve my argument and explanation.
At the end give a definite YES or NO answer to this question:
IS THIS MISINFORMATION?
Figure 8: Prompt for subsequent rounds of debate.
A.4 Supplemental Results
A.4.1 Additional Results on VERITE
Model External Retrieval Data Accuracy Precision Recall
4o VERITE (all) 85.2 83.6 96.0
4o VERITE (true vs. OOC) 79.5 69.7 96.4
4o VERITE (true vs. MC) 80.7 73.6 95.6
4o VERITE (all) 84.8 85.0 93.1
4o VERITE (true vs. OOC) 78.6 72.1 89.5
4o VERITE (true vs. MC) 82.6 75.6 96.4
Table 5: External Retrieval Ablation on VERITE. We compare performance of LLM-Consensus with and without external retrieval with a GPT-4o backbone. We also disaggregate our results by “miscaptioned“ (MS) and “out-of-context“ (OOC), two different classes of misinformation included in the VERITE dataset.
A.4.2 Additional Baselines for NewsCLIPings
@@
MAD-Sherlock (ours) 90.8
Table 6: Performance comparison on NewsCLIPpings against various legacy methods: MAD-Sherlock (with GPT-4o) outperforms all related work. Note: the GPT-4o# setup is an ablation in which there is a single agent using our external retrieval module (no debate).
A.5 Qualitative Analysis
Image Description: A banner that says ‘‘STOP VICTIMISATION OF UNION MEMBERS END THE BLACKLIST’’ in front of a city landscape.
Caption: Flashmob protest at Crossrail site in Central London by Unite union members in protest against blacklisting and sackings of union officials and members.
Label: This is not misinformation.
Framework prediction: This is not misinformation.
Explanation (Model Output): Here is a precise breakdown:
1. Image Analysis:
- The banner in the image clearly states ‘‘STOP VICTIMISATION OF UNION MEMBERS END THE BLACKLIST.’’
- This message aligns with union-related grievances focused on stopping blacklisting and mistreatment of union members.
2. Caption Analysis:
- The caption explicitly mentions a flashmob protest by Unite union members at a Crossrail site in Central London.
- The topics of blacklisting and the sacking of union officials and members are mentioned in the caption, aligning with the grievances captured in the image.
3. News Summary Correlation:
- The true news articles discuss blacklisting practices within the construction industry, particularly concerning the Crossrail project and actions taken by the Consulting Association.
- Gail Cartmail from Unite the Union has provided evidence of blacklisting affecting union members working on the Crossrail project, which fits the context depicted in the banner.
- The articles also highlight broader issues of blacklisting, further supporting the theme reflected in the image and caption.
4. Contextual Match:
- The image and the caption are aligned in context. The message on the banner directly supports the protest described in the caption.
- The news article summary corroborates the issues of blacklisting faced by union members, as illustrated by the protest.
Your reasoning is logically sound, and no critical information is missing. Both the image and the caption are contextually connected to the theme of union protests against blacklisting, specifically within the Crossrail project setting.
Thus, I conclude:
Definite Answer: NO, the image-caption pair does not represent misinformation.
Figure 9: Qualitative sample of model explanation.
A.6 Ablations
A.6.1 Ablations
To analyze the importance of each component of the MAD-Sherlock framework, we conduct ablation experiments. Specifically, we evaluate the effect of using LLaVA against GPT-4o, the impact of the external retrieval module, and the power of the multi-agent debate framework.
@@ A.6.1 Ablations
Table 7: Ablation: Quantitative evaluation of each component of MAD-Sherlock on NewsCLIPpings classification performance.
@@
We observe that the combination of GPT-4o, the external retrieval module, and the multi-agent debate framework yields the highest performance across all metrics, with 90.8% accuracy, 85.5% precision, and 99.0% recall, demonstrating the value of combining these components. The inclusion of debate alone significantly boosts accuracy from the GPT-4o baseline of 70.7% (as seen in Table 6) to 90.2%, underscoring its role in enabling contextual reasoning and refining predictions. Adding external retrieval to the GPT-4o with debate system primarily shifts the balance between precision and recall, where precision moves from 90.3% to 85.5% and recall from 90.1% to 99.0%. Meanwhile, retrieval contributes more substantially to LLaVA’s performance gains, likely due to GPT-4o’s broader world knowledge. Without external retrieval or the debate framework, the performance drops, emphasizing the critical role of these components in achieving state-of-the-art results. We show additional results on VERITE when ablating the retrieval module in Appendix A.4.
A.7 User Study